OpenStudy (anonymous):

Why is this expression called action \[ \int_0^t { (T - V) } dt\] where T is kinetic energy and V is potential energy.

OpenStudy (anonymous):

T-V = L (Lagrangian)

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Action_%28physics%29

OpenStudy (unklerhaukus):

\[[\text{energy}]\cdot[\text{time}]\]

OpenStudy (anonymous):

what exactly is action? how does [energy][time] give action?

OpenStudy (anonymous):

It's just a name, don't trouble yourself with the semantics of it.

OpenStudy (anonymous):

what is the significance of it?

OpenStudy (anonymous):

the Lagrange equation of motion is based on minimizing this value ...

OpenStudy (anonymous):

Yes. And when that's done, it reproduces newton's equations in cartesian coordinates.

OpenStudy (anonymous):

looks like i need to study calculus of variation before that.

OpenStudy (anonymous):

The calculus of variations can be used to derive the Lagrangian equations of motion, but is not necessary for actually using it.

OpenStudy (unklerhaukus):

a body will always take the part of least action

OpenStudy (anonymous):

I want to understand it ... must be something like Fermat's principle .. of least action

OpenStudy (unklerhaukus):

path*

OpenStudy (anonymous):

What exactly do you want to understand?

OpenStudy (anonymous):

that why \int T-V dt gives the trajectory

OpenStudy (anonymous):

If you minimize that function, L = T- V, then you get the following equation from the calculus of variations: \[ \frac{\partial L}{\partial x} = \frac{d}{dt} \frac{\partial L}{\partial \dot{x}}\] where \[\dot{x} = v_x \] If we assume T is independent of position (i.e. T = 1/2 mv^2) then the left hand side becomes \[\frac{-\partial U}{\partial x} \] And if U is independent of velocity (as is often the case) we can associate the right hand side with the momentum, since \[\frac{\partial T}{\partial \dot{x}} = m\dot{x} = p\] So we get \[\frac{d}{dt} p = - \frac{\partial U}{\partial x} = F\] which is just Newton's 2nd Law

OpenStudy (anonymous):

The Lagrangian formulation and the Newtonian formulation are equivalent statements, in the sense that you can derive one from the other, but the Lagrangian often has many advantages over the Newtonian picture. You need to start somewhere, and you can either start by saying "F = ma" or you can say that the Lagrangian is minimized and go from there.

OpenStudy (anonymous):

well ... i understand the application part. Just need to understand what the creators thought before formulating ... can we discuss this tomorrow? I need some time to work out. I got something on Google http://physics.stackexchange.com/questions/3912/does-action-in-classical-mechanics-have-a-interpretation

OpenStudy (anonymous):

also here http://physics.stackexchange.com/questions/9/hamiltons-principle

OpenStudy (anonymous):

They just wanted to reproduce newton's equations from a scalar that easily incorporated constraints, that's all.